OpenStudy (anonymous):

the x-intercept of the graph of f(x)=3log(x+5)-2 is A) 10^2/3 +5 B) 10^-2/3 +5 C) 10^-2/3 -5 D) 10^2/3 -5

OpenStudy (campbell_st):

well set f(x) =0 and solve so 0 = 3log(x + 5) -2 2 = 3log(x + 5) 2/3 = log(x + 5) \[10^{\frac{2}{3}} = x + 5\] it should be easy from here... hope it helps

OpenStudy (anonymous):

I got 10\[10^{\frac{ 2 }{ 3}} -5\]

OpenStudy (campbell_st):

that looks good to me